# A Comprehensive Survey of Direct Preference Optimization: Datasets, Theories, Variants, and Applications

### v2

@@ 3 Research Questions and Variants
Previous research derived pairwise preferences using pointwise rewards and the BT model. However, this approach was not comparable to direct pairwise preference modeling and failed to address inconsistencies within pairwise preferences. To overcome these limitations, some recent studies (Wu et al., 2024b; Rosset et al., 2024; Munos et al., 2024) have introduced Nash learning methodologies where each player is an LLM that outputs responses and aims to maximize its probability of being preferred over its opponent.
Inspired by generative adversarial networks (GAN) (Goodfellow et al., 2014), Self-Play fIne-tuNing (SPIN) considered a two-player game  (Wu et al., 2024b; Chen et al., 2024a), where the main player distinguishes the generated responses are from model or human, while the opponent player generates responses indistinguishable from human. This approach also eliminated the need for a reward model and derived a objective in a similar form to DPO.
An intuitive interpretation of DPO would lead one to believe it increases the likelihood of chosen responses while decreasing the likelihood of rejected responses. However, this does not support a well-observed phenomenon in which the likelihood of the chosen responses actually decreases over time (Pal et al., 2024; Rafailov et al., 2024b). To address such issues, some works focus  (Pang et al., 2024; Yuan et al., 2023; Xu et al., 2024b) on introducing extra objective terms.
Yuan et al. (2023) added a negative log-likelihood (NLL) loss similar to SFT (supervised fine-tuning) to rank loss. By doing so, They force the model to learn the response with the highest reward.  Xu et al. (2024b) incorporated a behavior cloning (BC) regularizer (Hejna et al., 2024) to ensure that $\pi_{\theta}$ does not deviate from the preferred data distribution. They prove that the regularize can boil down to adding a NLL term on the preferred data distribution.
@@
RQ3: Effect of KL Penalty Coefficient and Reference Model. As discussed in RQ1, DPO implicitly learns a reward model $r_{\theta}$ given an input $x$ and an output $r_{\theta}(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{ref}(y\mid x)}$, where $\pi_{\theta}$ and $\pi_{\mathrm{ref}}$ denote the distributions parameterized by the fine-tuned LLM and the reference LLM, respectively, and $\beta$ controls the strength of the Kullback-Leibler (KL) divergence regularization applied from the reference LLM. Some recent studies (Wu et al., 2024a; Liu et al., 2024c; Rafailov et al., 2024a; Xu et al., 2024b; Hong et al., 2024; Meng et al., 2024) have investigated the impact of the KL penalty coefficient $\mathbf{\beta}$ and the choice of reference model $\pi_{\mathrm{ref}}$.
@@
In KL-constrained RL and DPO, the KL penalty coefficient controls the trade-off between maximizing the reward and minimizing the deviation from the reference policy.  Wu et al. (2024a) introduces a novel framework that dynamically calibrates $\beta$ at the batch level, informed by data quality considerations. Additionally, this method incorporates $\beta$-guided data filtering to safeguard against the influence of outliers.  Liu et al. (2024c) have explored the optimal KL constraint strength for DPO, finding that a smaller KL constraint generally improves performance until the constraint becomes too small and leads to performance degradation. Empirically, following token log-probability difference experimental settings (Rafailov et al., 2024a), results reveal that as the strength of the KL constraint decreases, the DPO-fine-tuned model begins to assign significantly different probabilities to a small subset of specific tokens compared to the reference model.
@@
Given that the reference model requires double the memory and forward passes, recent research (Xu et al., 2024b; Meng et al., 2024) has explored alternative forms of regularization to replace the reference model. Besides, we usually initialize the reference model with a supervised fine-tuning (SFT) model, which involves a two-stage training process(i.e. SFT and preference optimization). Empirical evidence suggests that the effectiveness of DPO has strong reliance on the training effect of the LLMs after SFT (Feng et al., 2024). To address this issue, some studies integrate SFT into preference optimization (Hong et al., 2024).
@@
Xu et al. (2024b) argues that DPO (Rafailov et al., 2023) is memory-inefficient and speed-inefficient due to twice memory for reference and policy model and twice forward pass. To solve this issue, their proposed method Contrastive Preference Optimization (CPO) removes the reference model term by proving the upper boundary of DPO loss. Furthermore, CPO incorporates a behavior cloning (BC) regularize (i.e. negative log-likelihood NLL) to ensure that $\pi_{\theta}$ does not deviate from the preferred data distribution.
@@
Hong et al. (2024) observed that in SFT stage, the absence of penalty in Cross-Entropy loss hinders human preference learning. Inspired by unlikelihood penalty (Welleck et al., 2019), they incorporate an odds ratio-based penalty to negative log-likelihood (NLL), where $\text{odds}_{\theta}(y|x)=\frac{\pi_{\theta}(y|x)}{1-\pi_{\theta}(y|x)}$. Consequently, they proposed ORPO to incorporate SFT and preference optimization by the guidance of NLL and odds ratio (OR), where $\text{OR}_{\theta}(y_{w},y_{l})=\frac{\text{odds}_{\theta}(y_{w}|x)}{\text{odds}_{\theta}(y_{l}|x)}$, which requires neither an SFT warm-up stage nor a reference model. Besides, Compared to the probability ratio (PR), $\text{PR}_{\theta}(y_{w},y_{l})=\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}$, OR empirically avoids an overly extreme optimization to tokens in rejected response.
@@
RQ4: Online DPO. RLHF (Ouyang et al., 2022a) is an online framework for alignment, while recent advances in offline methods (e.g. DPO (Rafailov et al., 2023; Xu et al., 2024e)) show empirically efficient in practice. Online algorithms tend to be more computationally intensive than offline algorithms, due to sampling and training a reward model. Thus some studies have provided insights into the performance gap between online and offline algorithms.
@@
Yuan et al. (2024d) introduced the concept of Self-Rewarding to simultaneously enhance generation and reward performance, which consists of two stages: "Instruction Following" and "Self Rewarding". During the "Instruction Following" training, DPO was used to train the LLM to align with the instruction following dataset. In the "Self Rewarding" phase, candidate responses were evaluated by the model itself (LLM-as-a-judge (Zheng et al., 2023a)), and each candidate was assigned a score considering five metrics (i.e. relevance, coverage, usefulness, clarity, and expertise) to construct a preference dataset for preference optimization. For the next iteration, the fine-tuned model acts as the reference model. The two stages are iteratively performed until performance degradation.
Xu et al. (2024c) proposed Pairwise Cringe Optimization (PCO) to generalize Cringe Loss  (Adolphs et al., 2022) which is applied to binary feedback to pairwise feedback setting by using a soft margin extension. Besides, they also apply PCO iteratively in a way similar to  (Yuan et al., 2024d) outperforming DPO training in the same iterative way, while iterative PCO requires the reward mode to label generated responses.
Guo et al. (2024a) observed that LLMs can approximate well human labelling and can generate reliable preferences over responses, thus potentially could alleviate the distribution shift issue on reward models. Concurrently to Self-Rewarding (Yuan et al., 2024d),  Guo et al. (2024a) proposed an iterative DPO method called online AI feedback (OAIF) similar to Self-Rewarding. While Self-Rewarding obtains feedback from itself, OAIF can leverage feedback from any LLM, including stronger ones.
@@
Inspired by a phenomenon in nature that intraspecific competition drives species evolution,  Qi et al. (2024) proposed an Online Fast-Slow chasing DPO (OFS-DPO) for preference alignment. Specifically, they introduced two identical LoRA(Low-rank Adaptive) (Hu et al., 2021) modules with different optimization speeds using DPO with a new regularization term. Besides, in every certain step, they swap the roles of the fast module and slow module to maintain a more sustained gradient update momentum. This method has been validated by their theoretical analysis and empirical results.
@@
For instance, some recent research (Dubois et al., 2024; Singhal et al., 2023; Kabir et al., 2024) find that both humans and models tend to have a length exploitation whereby they prefer longer responses over shorter ones in pairwise preferences. For example, prior work  (Wang et al., 2023b) found that, given head-to-head model comparisons, there is a correlation of 0.96 between win rates and the average number of unique tokens in the model’s responses. Moreover, with such bias in preferences, aligned models tend to generate longer responses with potential hallucinations or repetitions.
@@
SimPO (Meng et al., 2024) modified the DPO loss to align the training objective with the generation objective, which also introduced a length normalization. Their empirical results demonstrated that length normalization increases the reward difference across all preference pairs, irrespective of their length. Additionally, they found that removing length normalization results in a strong positive correlation between reward and response length, leading to length exploitation.
@@
RQ6: Alignment Tax. RLHF aims to align models with human preferences. However, prior studies  (Ouyang et al., 2022a; Tu et al., 2023; OpenAI, 2024a; Anthropic, 2024; Wu et al., 2024c) have found a phenomenon known as the alignment tax. This phenomenon first refers to the improved preference accompanies degraded performances on other NLP tasks after alignment(Ouyang et al., 2022a).  Kim and Seo (2024) observed that the key neurons for safety and helpfulness significantly overlap, but they require different activation patterns of the shared neurons, which could account for the alignment tax. Therefore alignment tax poses a significant challenge to achieving effective and practical model alignment. Consequently, some studies (Lu et al., 2024a; Lin et al., 2024a; Fu et al., 2024; Lou et al., 2024a; Pentyala et al., 2024) have investigated the alignment tax and proposed methods to reduce its effect, focusing on data iteration, model merging (averaging).
@@
Compared to offline model merging methods  (Lin et al., 2024a),  Lu et al. (2024a) proposed an online method where at each optimization step of RLHF, they online merged the gradients of the policy model with the delta parameters of the SFT model. Their experimental findings indicate that this online approach is more effective than offline merging methods in reducing alignment tax while maximizing rewards.

### v3

@@ 3 Research Questions and Variants
Previous research derived pairwise preferences using pointwise rewards and the BT model. However, this approach was not comparable to direct pairwise preference modeling and failed to address inconsistencies within pairwise preferences. To overcome these limitations, some recent studies Wu et al. (2024b); Rosset et al. (2024); Munos et al. (2024) have introduced Nash learning methodologies where each player is an LLM that outputs responses and aims to maximize its probability of being preferred over its opponent.
Inspired by generative adversarial networks (GAN) Goodfellow et al. (2014), Self-Play fIne-tuNing (SPIN) considered a two-player game  Wu et al. (2024b); Chen et al. (2024a), where the main player distinguishes the generated responses are from model or human, while the opponent player generates responses indistinguishable from human. This approach also eliminated the need for a reward model and derived a objective in a similar form to DPO.
An intuitive interpretation of DPO would lead one to believe it increases the likelihood of chosen responses while decreasing the likelihood of rejected responses. However, this does not support a well-observed phenomenon in which the likelihood of the chosen responses actually decreases over time Pal et al. (2024); Rafailov et al. (2024b). To address such issues, some works focus  Pang et al. (2024); Yuan et al. (2023); Xu et al. (2024b) on introducing extra objective terms.
Yuan et al. (2023) added a negative log-likelihood (NLL) loss similar to SFT (supervised fine-tuning) to rank loss. By doing so, They force the model to learn the response with the highest reward.  Xu et al. (2024b) incorporated a behavior cloning (BC) regularizer Hejna et al. (2024) to ensure that $\pi_{\theta}$ does not deviate from the preferred data distribution. They prove that the regularize can boil down to adding a NLL term on the preferred data distribution.
@@
Lin et al. (2024d) proposed the cDPO algorithm, which improved reasoning capabilities by identifying and penalizing critical tokens. cDPO employed contrastive estimation to identify these negative-exclusive critical tokens efficiently. Departing from the example-level implicit reward mechanism of DPO, cDPO innovatively reframed it into a token-level explicit reward function, applying dynamic penalty weights to each token in negative trajectories. This systematically reduced the probability of generating highly critical tokens. Experiments demonstrated that cDPO significantly improved model accuracy on mathematical reasoning tasks.
Xiong et al. (2025) proposed Multi-turn Direct Preference Optimization (M-DPO) to address multi-turn mathematical reasoning tasks that integrate external tools. The approach formulates the problem as an MDP to theoretically re-parameterize the trajectory-level reward as a sum of log-probabilities over the agent’s actions. M-DPO implements this by "masking out" external observations during loss computation, forcing the model to focus solely on optimizing its own reasoning and execution steps. This prevents the model from learning to predict uncontrollable environmental messages.
@@
RQ3: Effect of KL Penalty Coefficient and Reference Model. As discussed in RQ1, DPO implicitly learns a reward model $r_{\theta}$ given an input $x$ and an output $r_{\theta}(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{ref}(y\mid x)}$, where $\pi_{\theta}$ and $\pi_{\mathrm{ref}}$ denote the distributions parameterized by the fine-tuned LLM and the reference LLM, respectively, and $\beta$ controls the strength of the Kullback-Leibler (KL) divergence regularization applied from the reference LLM. Some recent studies Wu et al. (2024a); Liu et al. (2024c); Rafailov et al. (2024a); Xu et al. (2024b); Hong et al. (2024); Meng et al. (2024) have investigated the impact of the KL penalty coefficient $\mathbf{\beta}$ and the choice of reference model $\pi_{\mathrm{ref}}$.
Figure 6: RQ3 Effect of KL Penalty Coefficient and Reference Model. The left panel illustrates the key role of the KL penalty in controlling the trade-off between reward maximization and policy deviation, showcasing both static (fixed coefficient) and dynamic (adaptive coefficient) strategies. The right panel categorizes methods based on their reliance on a reference model into two main classes: reference-based and reference-free approaches.
@@
In KL-constrained RL and DPO, the KL penalty coefficient controls the trade-off between maximizing the reward and minimizing the deviation from the reference policy.  Wu et al. (2024a) introduces a novel framework that dynamically calibrates $\beta$ at the batch level, informed by data quality considerations. Additionally, this method incorporates $\beta$-guided data filtering to safeguard against the influence of outliers.  Liu et al. (2024c) have explored the optimal KL constraint strength for DPO, finding that a smaller KL constraint generally improves performance until the constraint becomes too small and leads to performance degradation. Empirically, following token log-probability difference experimental settings Rafailov et al. (2024a), results reveal that as the strength of the KL constraint decreases, the DPO-fine-tuned model begins to assign significantly different probabilities to a small subset of specific tokens compared to the reference model.
@@
Given that the reference model requires double the memory and forward passes, recent research Xu et al. (2024b); Meng et al. (2024) has explored alternative forms of regularization to replace the reference model. Besides, we usually initialize the reference model with a supervised fine-tuning (SFT) model, which involves a two-stage training process(i.e. SFT and preference optimization). Empirical evidence suggests that the effectiveness of DPO has strong reliance on the training effect of the LLMs after SFT Feng et al. (2024). To address this issue, some studies integrate SFT into preference optimization Hong et al. (2024).
@@
Xu et al. (2024b) argues that DPO Rafailov et al. (2023) is memory-inefficient and speed-inefficient due to twice memory for reference and policy model and twice forward pass. To solve this issue, their proposed method Contrastive Preference Optimization (CPO) removes the reference model term by proving the upper boundary of DPO loss. Furthermore, CPO incorporates a behavior cloning (BC) regularize (i.e. negative log-likelihood NLL) to ensure that $\pi_{\theta}$ does not deviate from the preferred data distribution.
@@
Hong et al. (2024) observed that in SFT stage, the absence of penalty in Cross-Entropy loss hinders human preference learning. Inspired by unlikelihood penalty Welleck et al. (2019), they incorporate an odds ratio-based penalty to negative log-likelihood (NLL), where $\text{odds}_{\theta}(y|x)=\frac{\pi_{\theta}(y|x)}{1-\pi_{\theta}(y|x)}$. Consequently, they proposed ORPO to incorporate SFT and preference optimization by the guidance of NLL and odds ratio (OR), where $\text{OR}_{\theta}(y_{w},y_{l})=\frac{\text{odds}_{\theta}(y_{w}|x)}{\text{odds}_{\theta}(y_{l}|x)}$, which requires neither an SFT warm-up stage nor a reference model. Besides, Compared to the probability ratio (PR), $\text{PR}_{\theta}(y_{w},y_{l})=\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}$, OR empirically avoids an overly extreme optimization to tokens in rejected response.
@@
Gorbatovski et al. (2025) proposed a Trust Region DPO (TR-DPO) to address overoptimization in offline alignment by dynamically updating the reference policy during training. The method incorporates soft updates (interpolating the current policy with the reference policy) and hard updates (periodically replacing the reference policy) to reduce reliance on out-of-domain data, enabling higher policy deviation while maintaining generation quality. Experiments demonstrated that TR-DPO significantly outperforms standard DPO in dialogue and summarization tasks.
Lee et al. (2025) proposed $\epsilon$-DPO, which dynamically adjusted instance-level KL penalties through perturbation strategies. By observing the monotonicity of log-likelihood ratios between chosen and rejected responses under perturbed $\beta$ values, the algorithm adaptively determined KL coefficients without relying on batch-level statistics or reference policy updates.
Xie et al. (2025) proposed Exploratory Preference Optimization (XPO), an enhanced variant designed for online DPO settings. This work provides a profound insight that DPO implicitly performs a form of Q*-function approximation. Building on this observation, XPO ingeniously augments the standard DPO objective with an exploration bonus derived from the "optimism in the face of uncertainty" principle in reinforcement learning, encouraging the model to deliberately explore novel behaviors beyond the support of the initial reference policy. Through this simple modification, XPO theoretically overcomes the sample inefficiency of conventional online DPO caused by "passive exploration."
RQ4: Online DPO. RLHF Ouyang et al. (2022a) is an online framework for alignment, while recent advances in offline methods (e.g. DPO Rafailov et al. (2023); Xu et al. (2024e)) show empirically efficient in practice. Online algorithms tend to be more computationally intensive than offline algorithms, due to sampling and training a reward model. Thus some studies have provided insights into the performance gap between online and offline algorithms.
Figure 7: RQ4 Online DPO. This figure shows the training processes for Offline DPO and Online DPO. In the Online DPO, newly sample responses from each iteration can be annotated through various means, including by human annotators, the model itself, or any other stronger model.
@@
Yuan et al. (2024d) introduced the concept of Self-Rewarding to simultaneously enhance generation and reward performance, which consists of two stages: "Instruction Following" and "Self Rewarding". During the "Instruction Following" training, DPO was used to train the LLM to align with the instruction following dataset. In the "Self Rewarding" phase, candidate responses were evaluated by the model itself (LLM-as-a-judge Zheng et al. (2023a)), and each candidate was assigned a score considering five metrics (i.e. relevance, coverage, usefulness, clarity, and expertise) to construct a preference dataset for preference optimization. For the next iteration, the fine-tuned model acts as the reference model. The two stages are iteratively performed until performance degradation.
Xu et al. (2024c) proposed Pairwise Cringe Optimization (PCO) to generalize Cringe Loss  Adolphs et al. (2022) which is applied to binary feedback to pairwise feedback setting by using a soft margin extension. Besides, they also apply PCO iteratively in a way similar to  Yuan et al. (2024d) outperforming DPO training in the same iterative way, while iterative PCO requires the reward mode to label generated responses.
Guo et al. (2024a) observed that LLMs can approximate well human labelling and can generate reliable preferences over responses, thus potentially could alleviate the distribution shift issue on reward models. Concurrently to Self-Rewarding Yuan et al. (2024d),  Guo et al. (2024a) proposed an iterative DPO method called online AI feedback (OAIF) similar to Self-Rewarding. While Self-Rewarding obtains feedback from itself, OAIF can leverage feedback from any LLM, including stronger ones.
Figure 8: Data Curation and Iteration Strategies in Online DPO. The panel (a) shows a self-play framework that iteratively optimizes by pairing model responses against human responses. The panel (b) shows sDPO, which trains on data partitions with a dynamically updated reference model. The panel (c) shows the OPTune algorithm, which combines selective data regeneration with a weighted-DPO loss to accelerate optimization.
@@
Inspired by a phenomenon in nature that intraspecific competition drives species evolution,  Qi et al. (2024) proposed an Online Fast-Slow chasing DPO (OFS-DPO) for preference alignment. Specifically, they introduced two identical LoRA(Low-rank Adaptive) Hu et al. (2021) modules with different optimization speeds using DPO with a new regularization term. Besides, in every certain step, they swap the roles of the fast module and slow module to maintain a more sustained gradient update momentum. This method has been validated by their theoretical analysis and empirical results.
@@
Bose et al. (2025) proposed Hybrid Preference Optimization (HPO) to address the limitations of traditional DPO in sample efficiency and exploration capability by integrating offline preference data with online exploration mechanisms. The method theoretically established sample complexity bounds for policy optimization by relaxing concentrability conditions on offline data and introducing an optimistic regularizer for online feedback, which proved faster convergence rates than offline DPO or online exploration methods. HPO combined offline data and online exploration through an objective function that included a DPO loss on offline data and an exploration-aware regularizer, where parameter $\gamma$ balances offline-online contributions to approach optimal policies under limited samples.
@@
Figure 9: RQ5 Reward Hacking. Reward hacking occurs when a policy exploits the limited generalization of its reward model to achieve a high reward without meeting the actual objectives. A prominent example is length exploitation, where models generate longer responses to exploit data biases, often at the cost of quality. This figure shows various mitigation strategies to counter this issue, such as length normalization and regularization.
For instance, some recent research Dubois et al. (2024); Singhal et al. (2023); Kabir et al. (2024) find that both humans and models tend to have a length exploitation whereby they prefer longer responses over shorter ones in pairwise preferences. For example, prior work  Wang et al. (2023b) found that, given head-to-head model comparisons, there is a correlation of 0.96 between win rates and the average number of unique tokens in the model’s responses. Moreover, with such bias in preferences, aligned models tend to generate longer responses with potential hallucinations or repetitions.
@@
SimPO Meng et al. (2024) modified the DPO loss to align the training objective with the generation objective, which also introduced a length normalization. Their empirical results demonstrated that length normalization increases the reward difference across all preference pairs, irrespective of their length. Additionally, they found that removing length normalization results in a strong positive correlation between reward and response length, leading to length exploitation.
@@
Rashidinejad and Tian (2025) identified that conventional DPO suffers from two types of reward hacking: Type I arises from over-optimizing out-of-distribution actions due to partial data coverage, while Type II stems from performance degradation of the initial model when preference data poorly covers high-reward actions. They proposed POWER-DL, which combined robust reward optimization (POWER) with dynamic label weighting. POWER introduced weighted entropy regularization to enforce pessimism on uncovered actions, while dynamic labels adaptively suppressed destructive updates to the initial policy. This approach required only objective function modifications without additional model training or data augmentation.
RQ6: Alignment Tax. RLHF aims to align models with human preferences. However, prior studies  Ouyang et al. (2022a); Tu et al. (2023); OpenAI (2024a); Anthropic (2024); Wu et al. (2024c) have found a phenomenon known as the alignment tax. This phenomenon first refers to the improved preference accompanies degraded performances on other NLP tasks after alignmentOuyang et al. (2022a).  Kim and Seo (2024) observed that the key neurons for safety and helpfulness significantly overlap, but they require different activation patterns of the shared neurons, which could account for the alignment tax. Therefore alignment tax poses a significant challenge to achieving effective and practical model alignment. Consequently, some studies Lu et al. (2024a); Lin et al. (2024a); Fu et al. (2024); Lou et al. (2024a); Pentyala et al. (2024) have investigated the alignment tax and proposed methods to reduce its effect, focusing on data iteration, model merging (averaging).
Figure 10: RQ6 Alignment Tax. The scale in the left panel illustrates the "alignment tax"—the trade-off between preference accompanies and general capabilities. The right panel outlines various strategies to mitigate this issue, primarily focusing on data-level and model-level strategies.
@@
Compared to offline model merging methods  Lin et al. (2024a),  Lu et al. (2024a) proposed an online method where at each optimization step of RLHF, they online merged the gradients of the policy model with the delta parameters of the SFT model. Their experimental findings indicate that this online approach is more effective than offline merging methods in reducing alignment tax while maximizing rewards.
@@
Xiao et al. (2025) identified that DPO alignment introduced significant calibration errors in large language models, causing overconfidence in predictions and degrading reliability on tasks like TruthfulQA. They proposed RCFT, which integrated dynamic confidence adjustment mechanisms during calibration-aware fine-tuning and employed temperature scaling during inference to align model output probabilities with ground truth. This approach required only loss function modifications without additional data or constraints. Experiments demonstrated RCFT reduces calibration errors on models while maintaining DPO’s alignment performance.